OpenStudy (dr3am3r16):

i really need help with my firework portfolio in algebra 2

OpenStudy (lgnd_88):

alright

OpenStudy (lgnd_88):

Explain which part you need help with

OpenStudy (dr3am3r16):

all of it, math is not my strong subject at all, and i have no clue what im even supposed to do

OpenStudy (lgnd_88):

haha happy to help

OpenStudy (dr3am3r16):

awesome

OpenStudy (lgnd_88):

Okay so the initial eqaution is h=-16t^2+vt+s, that is from an object falling from a tall structure to the ground below and to get said equation is how fast the object fell and the amount of time it took to hit the ground

OpenStudy (lgnd_88):

take this example for an example haha h=-16t^2+vt+s, where h is the height (in feet), t is the time in motion (in seconds), v is the initial velocity (in feet per second), and s is the initial height(in feet) A gymnast dismounts the uneven parallel bars at a height of 8 feet with an initial upward velocity of 8 feet per second. Find the time (in seconds) it takes for the gymnast to reach the ground. . from the statement of the problem we have: v is 8 ft per sec s is 8 ft set h to zero (ground) and solve for t: h=-16t^2+vt+s 0=-16t^2+8t+8 dividing both sides by 8: 0=-2t^2+t+1

OpenStudy (lgnd_88):

multiplying both sides by -1: 0=2t^2-t-1 0=2t^2-2t+t-1 0=(2t^2-2t)+(t-1) 0=2t(t-1)+(t-1) 0=(t-1)(2t+1) t = {-1/2, 1} throw out the negative solution leaving: t = 1 second

OpenStudy (lgnd_88):

I think that should help with the whole thing the rest I can't help you with because it says it's an experiment hope it helps I will be here if you have anymore questions:)

OpenStudy (dr3am3r16):

okay thnx

OpenStudy (lgnd_88):

No problem